OpenStudy (anonymous):

Find the value of K that would make the limit exist. Find the limit. lim(x->0) (2x^3-6)/(x^k+3)

OpenStudy (anonymous):

I meant approaches inifinity

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \lim_{x\to\infty}\frac{2x^3-6}{x^k+3} \hspace{.33em}\\~\\ \normalsize \text{if i put k=3 then } \hspace{.33em}\\~\\ = \lim_{x\to\infty}\frac{2x^3-6}{x^3+3} \hspace{.33em}\\~\\ = \lim_{x\to\infty}\frac{2-\frac{6}{x^3}}{1+\frac{3}{x^3}} \hspace{.33em}\\~\\ =\frac{2-\frac{6}{\infty^3}}{1+\frac{3}{\infty^3}} \hspace{.33em}\\~\\ =\frac{2-0}{1+0} \hspace{.33em}\\~\\ =2 \hspace{.33em}\\~\\ \normalsize \text{so for k=3 the limit exists } \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Thank you very much, I understand it now.

OpenStudy (mathmath333):

yw

OpenStudy (anonymous):

In fact, any value of \(k\ge3\) would work.